# Recalibrating binary probabilistic classifiers

### v2

@@ Abstract
Recalibration of binary probabilistic classifiers to a target prior probability is an important task in areas like credit risk management. We analyse methods for recalibration from a distribution shift perspective. Distribution shift assumptions linked to the area under the curve (AUC) of a probabilistic classifier are found to be useful for the design of meaningful recalibration methods. Two new methods called parametric covariate shift with posterior drift (CSPD) and ROC-based quasi moment matching (QMM) are proposed and tested together with some other methods in an example setting. The outcomes of the test suggest that the QMM methods discussed in the paper can provide appropriately conservative results in evaluations with concave functions like for instance risk weights functions for credit risk.
Keywords:
Probabilistic classifier posterior probability prior probability calibration distribution shift dataset shift credit risk.
@@ 1 Introduction
Occasionally binary probabilistic classifiers are learned on a training dataset and then are applied to a test dataset which reflects a joint distribution of features and labels different than the distribution of the training dataset. Actually, quite often it is unknown how much the training and test distributions differ because for the instances in the test dataset only the features but not the labels can be observed. Hence, the feature training and test distributions might be different while the posterior probabilities are identical for the training and test datasets. This kind of dataset shift is called covariate shift and is rather benign in principle as it would not require any change of the probabilistic classifier (Storkey [21]). If however, another type of dataset shift other than covariate shift is incurred, applying the classifier learned on the training dataset to the instances in the test dataset without any changes risks to generate unreliable predictions of the labels.
@@
In this paper, we study the situation where indeed no labels are observed in the test dataset but where there exists an estimate of the proportion of positive labels, i.e. an estimate of the test prior probability of the positive class. The problem is then to recalibrate the probabilistic classifier learned on the training dataset such that the mean of the recalibrated classifier on the test dataset matches the estimate of the positive labels proportion.
@@
This is a common situation in credit risk management, see for instance Chapter 4 of Bohn and Stein [3]. So called probabilities of default (PDs) are estimated on a training sample with observations of solvent and defaulted borrowers and must be recalibrated before being evaluated for the borrowers in a live portfolio. The future solvency states of these borrowers in general are unknown but typically an estimate of the proportion of borrowers who are going to default is available. Sometimes, such estimates are conservative, i.e. they are likely to significantly overestimate the proportion of defaulters. Conservatism of the estimates can be a regulatory requirement or it could be part of stress testing exercises intended to assess the impact of unfavourable economic circumstances on the portfolio.
@@
Recalibration of posterior probabilities learned on a training dataset to a target prior probability on a test dataset is not a well-defined problem because there is more than one way to transform the original posterior probabilities such that the target is matched. Therefore, we are going to study the impact of the recalibration method on the values of concave or nearly concave functions of the posterior probabilities as an additional criterion to identify meaningful solutions. The risk weight functions for the calculation of minimum required capital under the “internal ratings based (IRB) approach” of the Basel credit risk framework (BCBS [2]) are a primary example of such functions. The findings of this paper can inform the choice of the recalibration method in credit risk management and similar contexts.
@@
This paper is organised as follows: Section 2 puts the paper into the context of related work. Section 3 describes the technical details of the setting of the paper and the recalibration problem. In addition, it introduces the evaluation of the solutions with a concave function as a criterion for assessing their appropriateness. Section 4 presents a number of methods for recalibration. Parametric CSPD (Section 4.4) and ROC-based QMM (Section 4.5) seem to be new. With an example in Section 5, we illustrate the dependence of the solutions to the recalibration problem on assumptions of distribution shift and identify some less reliable recalibration methods. Section 6 proposes a way forward to prudent recalibration and concludes the paper. Appendix 0.A provides additional technical details needed for the implementation of some of the methods discussed in Section 4.
2 Related work
@@
Calibration of probabilistic classifiers has often been treated in the literature, see the surveys by Ojeda et al. [14] and Silva Filho et al. [20]. In the typical calibration setting, a real-valued score is learned on a training dataset with joint observations of instances and labels. The score is then calibrated or mapped to become a probabilistic classifier on a test (or validation or calibration) dataset for which there are also joint observations of instances and labels. If the original score is already a probabilistic classifier then the term recalibration is sometimes used instead of calibration.
@@
Cautious calibration (Allikivi et al. [1]) is a variant of binary calibration with the goal to avoid either overconfidence or underconfidence of probabilistic classifiers. Like for calibration, it is assumed that the labels of the instances in the calibration and test datasets are known. A common feature with recalibration as defined in this paper is conservatism of the posterior estimates. In the setting of this paper, conservatism can be achieved by choosing an appropriate value for the target class-1 prior probability.
@@
Recalibration of probabilities of default (PDs) on a dataset without observation of labels but with knowledge of the prior probability of positive labels has been a topic of research for twenty or more years in credit risk (Bohn and Stein [3] and the references therein). Such recalibration may be considered an extreme case of learning with label proportions (Quadrianto et al. [16]) where there are no individual label observations but label proportions for groups of instances are available. In general, ‘learning with label proportions’ in the binary case requires that there are at least two groups of instances with different proportions of positive labels such that results from the related research are not applicable to the recalibration problem as studied in this paper.
@@
Quantification (or class distribution estimation, CDE) is another related problem. See Esuli et al. [8] for a recent survey. The goal of CDE in the binary case is to estimate the proportion of positive labels in a test dataset without any information on the labels. The primary common feature of the CDE and recalibration problems is the dependence of the solutions upon assumptions on the type of distribution shift between the training and test datasets. However, Remark 1 in Section 4.4 shows that the one-parameter version of the recalibration method ‘parametric CSPD’ may also be used for CDE.
@@
Covariate shift with posterior drift (CSPD) as introduced by Scott [19] turns out to be a useful assumption on the type of distribution shift for tackling the recalibration problem. See Sections 4.4 and 5 below.
@@ 3 Setting
In this paper, we consider binary, i.e. two-class classification problems for which we assume the following setting:
  • There are a class variable $Y$ with values in $\mathcal{Y}=\{0,1\}$ and a features (also called covariates) vector $X$ with values in $\mathcal{X}$. Each example (or instance) to be classified has a class label $Y$ and features $X$.
  • In the training dataset, for all examples their features $X$ and labels $Y$ are observed. $P$ denotes the training joint distribution, also called source distribution, of $(X,Y)$ from which the training dataset has been sampled.
  • In the test dataset, only the features $X$ of an example can immediately be observed. Its label $Y$ becomes known only with delay or not at all. $Q$ denotes the test joint distribution, also called target distribution, of $(X,Y)$ from which the test dataset has been sampled.
  • For the sake of a more concise notation, we write for short $p=P[Y=1]$ and $q=Q[Y=1]$ and assume $10<p<$ 1 and $10<q<$ 1.
[…]
The setting described above is called dataset shift (Storkey [21]) or distribution shift (Lipton et al. [12]) if source and target distribution of $(X,Y)$ are not the same, i.e. in case of $P(X,Y)\neq Q(X,Y)$.
@@ 3.1 Recalibration
We assume that since the joint source distribution $P(X,Y)$ of features $X\in\mathcal{X}$ and class $Y\in\{0,1\}=\mathcal{Y}$ is given, also the posterior probability $\eta_{P}(X)=P[Y=1\,|\,X]$ is known or can be estimated from the training dataset. $\eta_{P}(X)$ is a special case of probabilistic classifiers $\eta$ which are real-valued statistics with $0\leq\eta\leq 1$ and typically intended to approximate $\eta_{P}(X)$. Probabilistic classifiers for their part are special cases of scores (or scoring classifiers) which are real-valued statistics intended to provide a ranking of the instances in the dataset in the sense that a high score suggests a high likelihood that the instance has class label $1$ (positive label).
By assumption only the target marginal feature distribution $Q(X)$ is observed while the target joint distribution $Q(X,Y)$ is unknown. Nonetheless, for the recalibration problem, we also assume the target marginal label distribution, specified by $q=Q[Y=1]$ to be known. This is not in contradiction to $Q(X,Y)$ being unknown because in general the ensemble of marginal distributions does not uniquely determine the joint distribution. We will encounter an example for this phenomenon in Section 5 below.
@@
$q\ =\ E_{Q}[\eta_{Q}(X)]\ =\ E_{Q}\bigl{[}T(\eta_{P}(X))\bigr{)}].$ (1a)
In the following, we call this problem recalibration of the posterior probability $\eta_{P}(X)$ to a new prior probability $q$ of class $1$ under the target distribution.
@@
appears quite natural but actually is rather strong. Indeed, by Theorems 32.5 and 32.6 of Devroye et al. [7], (1b) is equivalent to $\eta_{P}(X)$ being sufficient for $X$ with respect to $Y$ under $Q$, i.e. $Q[Y=1\,|\,X]\ =\ Q[Y=1\,|\,\eta_{P}(X)]$. This sufficiency property may be interpreted as ‘the information provided by $\eta_{P}(X)$ about $Y$ is as good as the information by the whole set of features $X$ under the target distribution $Q$’.
@@ 3.2 Non-uniqueness of recalibration
The recalibration problem is not well-posed in the sense that its solution is not unique. Therefore, we study it in a context where underestimating $E_{Q}\bigl{[}C(\eta_{Q}(X))\bigr{]}$ for some fixed concave function $C:[0,1]\to\mathbb{R}$ ought to be avoided. In general, it holds that (by Jensen’s inequality and Lemma 1.2 of Lalley [11])
@@
for any random variable $0\leq Z\leq 1$ with $E_{Q}[Z]=q$, and $Z=\eta_{Q}(X)$ in particular. The maximum value $C(q)$ is taken for constant $Z=q$, the minimum value $(1-q)\,C(0)+q\,C(1)$ is realised for $Z$ with $Q[Z=1]=q=1-Q[Z=0]$.
However, assuming a distribution shift between source $P$ and target $Q$ which results in a constant posterior probability $\eta_{Q}(x)=q$ appears too restrictive in most real world environments. In Section 5 below, we demonstrate that assuming preservation of classification performance as measured by AUC (Area Under the Curve, see next section) between source $P$ and target $Q$ strikes a sensible note between too restrictive and too tolerant assumptions for the estimation of $E_{Q}\bigl{[}C(\eta_{Q}(X))\bigr{]}$ under the target features distribution.
@@ 3.3 Area Under the Curve (AUC)
AUC (Area Under the Curve111‘Curve’ refers to ROC (Receiver Operating Characteristic). See Fawcett [9] and the references therein for the most common definitions of ROC and AUC.) is a popular measure of performance of a binary classifier, i.e. AUC is considered an appropriate measure of the classifier’s ability to predict the true class label of an instance. See Chen et al. [5] for related comments. Since AUC plays an important role in some of the recalibration methods discussed in the following sections, we present here the population-level (in contrast to sample-based) representations of AUC that are used in this paper.
@@
If the score $S$ is assumed to be large for instances with high likelihood to have class $1$ and small for instances with high likelihood to have class 00, then the AUC for $S$ is defined as
@@
where $P^{\ast}$ denotes the product measure of $P_{1}$ and $P_{0}$, and $S_{1}$ and $S_{0}$ are the coordinate projections of the space $\boldsymbol{\mathcal{S}}\times\boldsymbol{\mathcal{S}}$ on which $P^{\ast}$ is defined. As a consequence, $S_{1}$ and $S_{0}$ are independent and $P^{\ast}[S_{y}\leq s]=P_{y}[S\leq s]$ for $y\in\{0,1\}$ and all $s\in\boldsymbol{\mathcal{S}}$.
By Definition (3a), on the one hand $AUC_{S}$ is “equivalent to the probability that the classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance” (p. 868 of Fawcett [9]) if both of the class-conditional distribution functions of the score $S$ are continuous. On the other hand, it holds that $AUC_{S}=\frac{1}{2}$ for any uninformative score $S$ – i.e. in the case $P_{0}=P_{1}$ – even if both of the class-conditional score distribution functions have discontinuities. Note that computing $AUC_{S}$ by means of (3a) at first glance requires knowledge of the joint distribution $P(S,Y)$ of $S$ and $Y$ which would have to be inferred from a sample of paired $(S,Y)$ observations.
@@
To indicate the way $AUC_{S}$ is computed, in this paper we refer simply to AUC if $AUC_{S}$ is assumed to be computed or inferred by means of (3a) irrespectively of the origin of $P_{0}$ and $P_{1}$. We refer to implied AUC if $AUC_{S}$ is assumed to be computed by means of (3b) in combination with (3a). See Appendix 0.A.1 for a more detailed formula for implied AUC in the case of a discrete-valued score $S$.

### v3

@@ Abstract
Recalibration of binary probabilistic classifiers to a target prior probability is an important task in areas like credit risk management. However, recalibration of a classifier learned on a training dataset to a target on a test dataset in general is not a well-defined problem because there might be more than one way to transform the original posterior probabilities such that the target is matched. In this paper, methods for recalibration are analysed from a distribution shift perspective. Distribution shift assumptions linked to the area under the curve (AUC) of a probabilistic classifier are found to be useful for the design of meaningful recalibration methods. Two new methods called parametric covariate shift with posterior drift (CSPD) and ROC-based quasi moment matching (QMM) are proposed and tested together with some other methods in an example setting. The outcomes of the test suggest that the QMM methods discussed in the paper can provide appropriately conservative results in evaluations with concave functions like for instance risk weights functions for credit risk.
Keywords: Probabilistic classifier, calibration, distribution shift, dataset shift, credit risk.
MSC classes: 68T09, 91G40
@@ 1 Introduction
Occasionally binary probabilistic classifiers are learned on a training dataset (typically a collection of data points observed in the past) and then are applied to a test dataset which reflects a joint distribution of features and labels different than the distribution of the training dataset. Actually, quite often it is unknown how much the training and test distributions differ because for the instances in the test dataset only the features but not the labels can be observed. Hence, the feature training and test distributions might be different while the posterior probabilities are identical for the training and test datasets. This kind of dataset shift is called covariate shift and is rather benign in principle as it would not require any change of the probabilistic classifier (Storkey, 2009). If however another type of dataset shift other than covariate shift is incurred, applying the classifier learned on the training dataset to the instances in the test dataset without any change risks to generate unreliable predictions of the labels.
@@
This paper aims to study the situation where indeed no labels are observed in the test dataset but where there exists an estimate of the proportion of positive labels, i.e. an estimate of the test prior probability of the positive class. The problem is then to recalibrate the probabilistic classifier learned on the training dataset such that the mean of the recalibrated classifier on the test dataset matches the estimate of the proportion of positive labels.
@@
This is a common situation in credit risk management, see for instance Chapter 4 of Bohn and Stein (2009). So called probabilities of default (PDs) are estimated on a training sample with observations of solvent and defaulted borrowers and must be recalibrated before being evaluated for the borrowers in a live portfolio. The future solvency states of these borrowers in general are unknown but typically an estimate of the proportion of borrowers who are going to default is available. Sometimes, such estimates are conservative, i.e. they are likely to significantly overestimate the proportion of defaulters. Conservatism of the estimates can be a regulatory requirement or it could be part of stress testing exercises intended to assess the impact of unfavourable economic circumstances on the portfolio.
@@
Recalibration of posterior probabilities learned on a training dataset to a target prior probability on a test dataset is not a well-defined problem because there is more than one way to transform the original posterior probabilities such that the target is matched. Therefore, as an additional criterion to identify meaningful solutions the impact of the recalibration method on the values of concave or nearly concave functions of the posterior probabilities is investigated. The risk weight functions for the calculation of minimum required capital under the “internal ratings based (IRB) approach” of the Basel credit risk framework (BCBS, ) are a primary example of such functions. The findings of this paper can inform the choice of the recalibration method in credit risk management and similar contexts.
@@
This paper is organised as follows: Section 2 puts the paper into the context of related work. Section 3 describes the technical details of the setting of the paper and the recalibration problem. In addition, it introduces the evaluation of the solutions with a concave function as a criterion for assessing their appropriateness. Section 4 presents a number of methods for recalibration. Parametric CSPD (Section 4.4) and ROC-based QMM (Section 4.5) seem to be new. An example in Section 5 illustrates the dependence of the solutions to the recalibration problem on assumptions of distribution shift and helps to identify some less reliable recalibration methods. Section 6 proposes a way forward to prudent recalibration and concludes the paper. Appendix A provides additional technical details needed for the implementation of some of the methods discussed in Section 4.
2 Literature review
@@
Calibration of probabilistic classifiers has often been treated in the literature, see the surveys by Ojeda et al. (2023) and Silva Filho et al. (2023). In the typical calibration setting, a real-valued score is learned on a training dataset with joint observations of instances and labels. The score is then calibrated or mapped to become a probabilistic classifier on a test (or validation or calibration) dataset for which there are also joint observations of instances and labels. If the original score is already a probabilistic classifier then the term recalibration is sometimes used instead of calibration. Recent work by Moreo (2025) shows that, under distribution shift, the problem of classifier calibration is closely related to the problems of quantifying prior class probabilities and predicting the accuracy of a classifier.
@@
Cautious calibration (Allikivi et al., 2024) is a variant of binary calibration with the goal to avoid either overconfidence or underconfidence of probabilistic classifiers. Like for calibration, it is assumed that the labels of the instances in the calibration and test datasets are known. A common feature with recalibration as defined in this paper is conservatism of the posterior estimates. In the setting of this paper, conservatism can be achieved by choosing an appropriate value for the target class-1 prior probability.
@@
Recalibration of probabilities of default (PDs) on a dataset without observation of labels but with knowledge of the prior probability of positive labels has been a topic of research for twenty or more years in credit risk (Bohn and Stein, 2009, and the references therein). Such recalibration may be considered an extreme case of learning with label proportions (Quadrianto et al., 2009) where there are no individual label observations but label proportions for groups of instances are available. In general, ‘learning with label proportions’ in the binary case requires that there are at least two groups of instances with different proportions of positive labels such that results from the related research are not applicable to the recalibration problem as studied in this paper.
@@
Quantification (or class distribution estimation, CDE) is another related problem. See Esuli et al. (2023) for a recent survey. The goal of CDE in the binary case is to estimate the proportion of positive labels in a test dataset without any information on the labels. The primary common feature of the CDE and recalibration problems is the dependence of the solutions upon assumptions on the type of distribution shift between the training and test datasets. However, Remark 4.2 in Section 4.4 shows that the one-parameter version of the recalibration method ‘parametric CSPD’ may also be used for CDE.
@@
Covariate shift with posterior drift (CSPD) as introduced by Scott (2019) turns out to be a useful assumption on the type of distribution shift for tackling the recalibration problem. See Sections 4.4 and 5 below.
@@ 3 Setting
In this paper, the focus is on the binary case, i.e. two-class classification problems for which the following setting is assumed:
  • There are a class variable $Y$ with values in some measurable space $\mathcal{Y}=\{0,1\}$ and a features (also called covariates) vector $X$ with values in $\mathcal{X}$. Each example (or instance) to be classified has a class label $Y$ and features $X$.
  • In the training dataset, for all examples their features $X$ and labels $Y$ are observed. $P$ denotes the training joint distribution, also called source distribution of $(X,Y)$ from which the training dataset has been sampled.
  • In the test dataset, only the features $X$ of an example can immediately be observed. Its label $Y$ becomes known only with delay or not at all. $Q$ denotes the test joint distribution, also called target distribution of $(X,Y)$ from which the test dataset has been sampled.
  • For the sake of a more concise notation, write for short $p=P[Y=1]$ and $q=Q[Y=1]$ and assume $0<p<1$ and $0<q<1$.
In addition, the notation $E_{P}[Z]=\int Z\,dP$ and $E_{Q}[Z]=\int Z\,dQ$ is used for integrable real-valued random variables $Z$.
The setting described above is called dataset shift (Storkey, 2009) or distribution shift (Lipton et al., 2018) if source and target distribution of $(X,Y)$ are not the same, i.e. in case of $P(X,Y)\neq Q(X,Y)$.
@@ 3.1 Recalibration
Assume that since the joint source distribution $P(X,Y)$ of features $X\in\mathcal{X}$ and class $Y\in\{0,1\}=\mathcal{Y}$ is given, also the posterior probability $\eta_{P}(X)=P[Y=1\,|\,X]$ is known or can be estimated from the training dataset. $\eta_{P}(X)$ is a special case of probabilistic classifiers $\eta$ which are real-valued statistics with $0\leq\eta\leq 1$ and typically intended to approximate $\eta_{P}(X)$. Probabilistic classifiers for their part are special cases of scores (or scoring classifiers) which are real-valued statistics intended to provide a ranking of the instances in the dataset in the sense that a high score suggests a high likelihood that the instance has class label $1$ (positive label).
By assumption only the target marginal feature distribution $Q(X)$ is observed while the target joint distribution $Q(X,Y)$ is unknown. Nonetheless, for the recalibration problem, it is assumed that the target marginal label distribution, specified by $q=Q[Y=1]$ to be known. This is not in contradiction to $Q(X,Y)$ being unknown because in general the ensemble of marginal distributions does not uniquely determine the joint distribution. An example for this phenomenon is presented in Section 5 below.
@@
$q\ =\ E_{Q}[\eta_{Q}(X)]\ =\ E_{Q}\bigl[T(\eta_{P}(X))\bigr].$ (1a)
In the following, this problem is called recalibration of the posterior probability $\eta_{P}(X)$ to a new prior probability $q$ of class $1$ under the target distribution.
@@
appears quite natural but actually is rather strong. Indeed, by Theorems 32.5 and 32.6 of Devroye et al. (1996), (1b) is equivalent to $\eta_{P}(X)$ being sufficient for $X$ with respect to $Y$ under $Q$, i.e. $Q[Y=1\,|\,X]\ =\ Q[Y=1\,|\,\eta_{P}(X)]$. This sufficiency property may be interpreted as ‘the information provided by $\eta_{P}(X)$ about $Y$ is as good as the information from the whole set of features $X$ under the target distribution $Q$’.
@@ 3.2 Non-uniqueness of recalibration
The recalibration problem is not well-posed in the sense that its solution is not unique. Therefore, a context is assumed where underestimating $E_{Q}\bigl[C(\eta_{Q}(X))\bigr]$ for some fixed concave function $C:[0,1]\to\mathbb{R}$ ought to be avoided. In general, it holds that (by Jensen’s inequality and Lemma 1.2 of Lalley, 2013)
@@
for any random variable $0\leq Z\leq 1$ with $E_{Q}[Z]=q$, and $Z=\eta_{Q}(X)$ in particular. The maximum value $C(q)$ is taken for constant $Z=q$, the minimum value (1q)C(0)+qC(1)(1-q)\,C(0)+q\,C(1) is realised for $Z$ with Q[Z=1]=q=1Q[Z=0]Q[Z=1]=q=1-Q[Z=0].
However, assuming a distribution shift between source $P$ and target $Q$ which results in a constant posterior probability $\eta_{Q}(x)=q$ appears too restrictive in most real world environments. In Section 5 below, it is demonstrated that assuming preservation of classification performance as measured by AUC (Area Under the Curve, see next section) between source $P$ and target $Q$ strikes a sensible note between too restrictive and too tolerant assumptions for the estimation of $E_{Q}\bigl[C(\eta_{Q}(X))\bigr]$ under the target features distribution.
@@ 3.3 Area Under the Curve (AUC)
AUC (Area Under the Curve111‘Curve’ refers to ROC (Receiver Operating Characteristic). See Fawcett (2006) and the references therein for the most common definitions of ROC and AUC.) is a popular measure of performance of a binary classifier, i.e. AUC is considered an appropriate measure of the classifier’s ability to predict the true class label of an instance. See Chen et al. (2018) for related comments. Since AUC plays an important role in some of the recalibration methods discussed in the following sections, in this section population-level (in contrast to sample-based) representations of AUC are presented for use in the following.
@@
If the score $S$ is assumed to be large for instances with high likelihood to have class $1$ and small for instances with high likelihood to have class 0, then the AUC for $S$ is defined as
@@
where $P^{\ast}=P_{1}\otimes P_{0}$ denotes the product measure of $P_{1}$ and $P_{0}$, and $S_{1}$ and $S_{0}$ are the coordinate projections of the space $\boldsymbol{\mathcal{S}}\times\boldsymbol{\mathcal{S}}$ on which $P^{\ast}$ is defined. As a consequence, $S_{1}$ and $S_{0}$ are independent and $P^{\ast}[S_{y}\leq s]=P_{y}[S\leq s]$ for $y\in\{0,1\}$ and all $s\in\boldsymbol{\mathcal{S}}$.
By Definition (3a), on the one hand $AUC_{S}$ is “equivalent to the probability that the classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance” (Fawcett, 2006, pp. 868) if both of the class-conditional distribution functions of the score $S$ are continuous. On the other hand, it holds that $AUC_{S}=\frac{1}{2}$ for any uninformative score $S$ – i.e. in the case $P_{0}=P_{1}$ – even if both of the class-conditional score distribution functions have discontinuities. Note that computing $AUC_{S}$ by means of (3a) at first glance requires knowledge of the joint distribution $P(S,Y)$ of $S$ and $Y$ which would have to be inferred from a sample of paired $(S,Y)$ observations.
@@
To indicate the way $AUC_{S}$ is computed, in this paper, the notion of AUC is used if $AUC_{S}$ is assumed to be computed or inferred by means of (3a) irrespectively of the origin of $P_{0}$ and $P_{1}$. A reference to implied AUC is used if $AUC_{S}$ is assumed to be computed by means of (3b) in combination with (3a). See Appendix A.1 for a more detailed formula for implied AUC in the case of a discrete-valued score $S$.